# SCAIL: Towards Studio-Grade Character Animation via In-Context Learning of 3D-Consistent Pose Representations

### v2

@@ 5 Conclusion
Performance Tradeoff. For the injection of the representation in full-context settings, we observe that the rendered pose sequence is relatively sparse as most frame areas consist of non-informative black pixels. In our method, spatial pooling to pose sequence can serve as a simple workaround to effectively reduce the contextual pose tokens to $1/4$ and preserves accurate pose-following capability. In addition, full-context pose injection introduces no new parameters except the additional patchify layer to the original model, offering a more streamlined architecture compared to stacking additional DiT layers in residual context-tuning methods [17]. In general, considering the significant motion error reduction, we conclude that the modest efficiency trade-off is acceptable, particularly in studio-grade scenarios which prioritize stringent accuracy and stability.
Methods Self-Driven Animation
[…]
SCAIL-1.3B (Ours) 18.08 0.639 0.249 228.62
SCAIL-1.3B w/ 2D Pose 17.08 0.619 0.284 295.36
Table A1: Ablation studies to validate the 3D pose representation on the self-driven subset of Studio-Bench.
@@
A2.1 Ablation on the Pose Representation
To validate the effectiveness of our proposed 3D-consistent pose representation, we compare with a 2D keypoints-based pipeline (extracted using DWPose [46] and denoted as w/ 2D Pose). Specifically, we implement the 2D augmentation strategy as close as to the 3D version, with similar figure settings and same augmentation ratio of 0.8. Faces and hands are also augmented identically to the 3D version and retargeting logic from [38] are applied during inference. All model architectures keep the same using full-context pose injection and P-RoPE. As shown in Table A1, the estimation noise and the 2D-3D mismatch noise introduced by the pipeline significantly undermine the model’s performance in the self-driven subset of our challenging Studio-Bench, which involves a high ratio of complex motions.
@@
Methods Self-Driven Animation
[…]
SCAIL-1.3B (Ours) 18.08 0.639 0.249 228.62
w/o augmentation 18.09 0.635 0.251 230.31
Table A2: Ablation studies to validate the augmentation strategy on the self-driven subset of Studio-Bench.
@@
A3 Details on Evaluation
A3.1 Details on Studio-Bench
@@
A3.2 More Examples on Studio-Bench
@@
A4 Discussion
A4.1 Limitations and Future Work

### v3

@@ 5 Conclusion
Performance Tradeoff. For the injection of the representation in full-context settings, we observe that the rendered pose sequence is relatively sparse as most frame areas consist of non-informative black pixels. In our method, spatial pooling to pose sequence can serve as a simple workaround to effectively reduce the contextual pose tokens to $1/4$ and preserves accurate pose-following capability. In addition, full-context pose injection introduces no new parameters except the additional patchify layer to the original model, offering a more streamlined architecture compared to stacking additional DiT layers in residual context-tuning methods [17]. For quantitative comparison, we compare the inference costs of the two injection schemes (512 * 896 resolution, 81 frames, 20 diffusion steps) on an H100 GPU. In general, considering the significant motion error reduction, we conclude that the modest efficiency trade-off is acceptable, particularly in studio-grade scenarios which prioritize stringent accuracy and stability.
Methods (all w/ CFG) Inference Time (s) FPS Memory (GB)
14B w/ channel-concat 286.11 0.283 61.7
14B w/ full-context (Ours) 380.78 (+33.1%) 0.213 (-24.9%) 68.5 (+11.0%)
@@
A2.1 Ablation details on the Pose Representation
As mentioned in the main paper, we implement the 2D augmentation strategy as close as to the 3D version, with similar figure settings and same augmentation ratio of 0.8. Faces and hands are also augmented identically to the 3D version and retargeting logic from [38] are applied during inference. As shown in Table 2, the estimation noise and the 2D-3D mismatch noise introduced by the pipeline significantly undermine the model’s performance in the self-driven subset of our challenging Studio-Bench, which involves a high ratio of complex motions.
@@
A3 Data Source
Our training dataset is composed of three primary data sources: (1) samples retrieved from our internal base model training data and other downstream tasks, (2) a large collection of high-resolution dance videos from Bilibili and YouTube, and (3) additional sports videos such as gymnastics and figure skating. To ensure diversity, we maintain a certain proportion of stylized content, including 3D and 2D animations from source (1), as well as MMD and Live2D animations from source (2).
A4 Details on Studio-Bench
@@
A4.1 More Examples on Studio-Bench
@@
A5 Discussion
A5.1 Limitations and Future Work